OpenStudy (anonymous):

Please help me figure out how to find the answer? cos(x+y) if tanx=5/3 and siny=1/3 [answer: (12sqrt17-5sqrt34)/102

OpenStudy (anonymous):

I got \[cosy=(2\sqrt3)/2\] and \[tany=\sqrt3/6\] but idk if those are right.

OpenStudy (anonymous):

I don't know how to find cosx and sinx from there.

OpenStudy (jdoe0001):

first off.... I assume you know the sum identity for cos(x+y) ?

OpenStudy (anonymous):

Yes, \[\cos(x+y)=cosxcosy \pm sinxsiny\]

OpenStudy (jdoe0001):

right

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

\(\bf tan(x)=\cfrac{5}{3}\implies \cfrac{opposite}{adjacent}\implies \cfrac{b}{a}\implies \cfrac{b=5}{a=3}\\ \quad \\ cos(y)=\cfrac{a}{{\color{blue}{ c}}}\qquad c^2=a^2+b^2\implies {\color{blue}{ c}}=\sqrt{a^2+b^2}\\ \quad \\ sin(y)=\cfrac{b}{{\color{blue}{ c}}}\\ \quad \\ ------------------------\\ sin(y)=\cfrac{1}{3}\implies \cfrac{opposite}{hypotenuse}\implies \cfrac{b}{c}\implies \cfrac{b=1}{c=3}\\ \quad \\ cos(x)=\cfrac{{\color{blue}{ a}}}{c}\qquad c^2=a^2+b^2\implies \sqrt{c^2-b^2}={\color{blue}{ a}}\)

OpenStudy (jdoe0001):

hmmm shoot, I got my haemm angles a bit.. anyhow \(\bf tan(x)=\cfrac{5}{3}\implies \cfrac{opposite}{adjacent}\implies \cfrac{b}{a}\implies \cfrac{b=5}{a=3}\\ \quad \\ cos(x)=\cfrac{a}{{\color{blue}{ c}}}\qquad c^2=a^2+b^2\implies {\color{blue}{ c}}=\sqrt{a^2+b^2}\\ \quad \\ sin(x)=\cfrac{b}{{\color{blue}{ c}}}\\ \quad \\ ------------------------\\ sin(y)=\cfrac{1}{3}\implies \cfrac{opposite}{hypotenuse}\implies \cfrac{b}{c}\implies \cfrac{b=1}{c=3}\\ \quad \\ cos(y)=\cfrac{{\color{blue}{ a}}}{c}\qquad c^2=a^2+b^2\implies \sqrt{c^2-b^2}={\color{blue}{ a}}\)

OpenStudy (jdoe0001):

once you get the missing, using the pythagorean theorem, you'd know the sine and cosine for angles "x" and "y" once you have those, just plug them in in the cos(x+y) sum identity

OpenStudy (anonymous):

I got the answer!! Thank you.

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Would you mind doing one more? sec(x-y) if cscx=5/3 and tany=12/5

OpenStudy (jdoe0001):

tis the same... really, just using the pythagorean theorem to find the missing side(s) \(\bf csc(x)=\cfrac{5}{3}\implies \cfrac{hypotenuse}{opposite}\implies \cfrac{c}{b}\implies \cfrac{c=5}{b=3}\\ \quad \\ tan(y)=\cfrac{12}{5}\implies \cfrac{opposite}{adjacent}\implies \cfrac{b}{a}\implies \cfrac{b=12}{a=5}\)

OpenStudy (jdoe0001):

just keep in mind that \(\bf sec(\theta)=\cfrac{1}{cos(\theta)}\qquad \qquad sec(x-y)=\cfrac{1}{cos(x-y)}\)